Based on your knowledge of the flexible accelerator model of investment, suppose you have information on initial capital stock in year 2009 to be Ȼ10,000, and that desired capital stock is Ȼ80,000. If the factor by which firms plan to close the gap between desired capital stock and actual capital stock is 0.15, how many years will take for the firms to meet the desired capital stock?
1
Expert's answer
2020-06-02T09:01:20-0400

flexible accelerator model of investment:

"(80000-10 000)\times0.15=10 500"

"10000+10500\times x\times0.15=80 000"

"1575\times x=70 000"

x=44.44


44  years